# Quality-Aware Image-Text Alignment for Real-World Image Quality Assessment

### v2

@@ Acknowledgments
We compare the robustness of QualiCLIP with CLIP-IQA by conducting the group maximum differentiation (gMAD) competition [22]. We provide more details on gMAD in Sec. 4.3. Fig. S1 shows the results. When QualiCLIP is fixed (Fig. 1(a)), CLIP-IQA struggles to identify image pairs with an evident quality gap. In contrast, when QualiCLIP operates as the attacker (Fig. 1(b)), it successfully highlights the failures of CLIP-IQA by finding image pairs displaying significantly different quality. This result shows that our approach demonstrates greater robustness than CLIP-IQA.
@@
S2.4 t-SNE Visualization
@@
S2.5 Inference Time
@@
To validate this, we compare the average inference time of our model with that of the supervised baselines on the KonIQ-10k dataset [12], which comprises 10K $1024\times 768$px images. We conduct the experiments on an NVIDIA RTX 2080Ti GPU. We report the results in Tab. S6. As expected, QualiCLIP has a similar inference time to models based solely on ResNet50, such as CLIP-IQA+, LIQE, CONTRIQUE, and ARNIQA. Note that CONTRIQUE and ARNIQA employ the full- and half-scale input image to compute the final quality score, and thus are more computationally expensive. Moreover, our model is faster than methods based on two encoders, namely GRepQ and Re-IQA. Notably, Re-IQA, like QualiCLIP, is designed for real-world IQA. This outcome demonstrates the efficiency and applicability of our approach in real-world scenarios.
@@
Table S6: Comparison of the average inference time between QualiCLIP and supervised approaches on the KonIQ-10k dataset.
@@
Training details
We rely on a ResNet50 [11] as the backbone for CLIP’s image encoder. Similar to [41], we remove the positional embedding from the encoder to make our model capable of taking images of any resolution as input. The dimension $d$ of CLIP’s embedding space is 1024. Differently from [36], we do not train a projector head on top of CLIP’s image encoder. We keep CLIP’s text encoder frozen. We train our model for 10 epochs and we use early stopping based on validation performance to reduce the overall training time. We employ an AdamW [19] optimizer with a weight decay and a learning rate of $21e\!-\!$ 2 and $91e\!-\!$ 9, respectively. During training, we use a patch size of 224 and a batch size of 16. We set the margins $m_{cons}$ in Eq. 1 and $m_{rank}$ in Eqs. 2 and 3 to $32.5e\!-\!$ 3 and $26.75e\!-\!$ 2, respectively. The loss weights $\lambda_{cons}$, $\lambda_{pos}$ and $\lambda_{neg}$ in Eq. 4 are all equal to 1. We set the temperature hyperparameter $\tau$ in Eq. 5 to 2. At inference time, our model takes the whole image as input.
Prompts
Following [42, 43], we employ multiple pairs of antonym prompts during training and inference. In particular, we use: 1) “Good/Bad photo”; 2) “Good/Bad picture”; 3) “High-resolution/Low-resolution image”; 4) “High-quality/Low-quality image”; 5) “Sharp/Blurry image”; 6) “Sharp/Blurry edges”; 7) “Noise-free/Noisy image”. We average the similarities between the images and the pairs of prompts.
Synthetic distortions
As detailed in Sec. 3.2, during training we synthetically degrade pristine images with increasing intensity levels. Specifically, similar to [1], we consider 24 distinct distortion types divided into the 7 degradation groups defined by the KADID-10k [17] dataset. Each degradation has 5 degrees of progressively higher intensity. We report an example for all the intensity levels of each type of degradation in Figs. S3, S4, S5, S6, S7, S8 and S9. Each distortion is described as follows:

### v3

@@ Acknowledgments
To assess the robustness of our model, we carry out the group maximum differentiation (gMAD) competition [24]. In particular, we compare QualiCLIP against GRepQ and CLIP-IQA using the Waterloo Exploration Database [23] dataset, which comprises 95K synthetically degraded images without MOS annotations. In this evaluation, one model is fixed to function as a defender, and its quality predictions are grouped into two distinct levels. The other model assumes the role of the attacker, tasked with identifying image pairs within each level that exhibit the greatest quality difference. For a model to demonstrate robustness, the selected image pairs should show comparable quality when acting as the defender while exhibiting a notable quality disparity when assuming the role of the attacker. We observe that when we fix QualiCLIP at a low-quality level (Fig. 1(a) top), GRepQ fails to find picture pairs with an obvious quality difference. When considering a high-quality level (Fig. 1(a) bottom), the image pair identified by GRepQ shows a slight quality gap. However, when assuming the role of the attacker (Fig. 1(b)), QualiCLIP successfully exposes the failures of GRepQ, as it pinpoints image pairs displaying a significant quality disparity. Fig. S2 shows that the same considerations can be drawn when analyzing the results of the gMAD competition between QualiCLIP and CLIP-IQA. Hence, our approach demonstrates greater robustness than GRepQ and CLIP-IQA.
Figure S1: gMAD competition results between QualiCLIP and GRepQ [41]. (a): Fixed QualiCLIP at a low- (top) and high-quality (bottom) level, respectively. (b): Fixed GRepQ at a low- (top) and high-quality (bottom) level, respectively.
@@
(a) Positive prompt
(b) Negative prompt
Figure S3: gradCAM visualization of the most important regions of the input image for each of the antonym prompts.
S2.4 gradCAM visualization
We evaluate the explainability of our model and CLIP-IQA via a gradCAM [36] visualization. gradCAM is a visualization technique aimed at understanding which regions of an input image are most influential for a model’s decision by studying the gradients of a given layer. We employ gradCAM to produce a heatmap of the regions of the image that activate the most for each of the antonym prompts. We employ “Good photo” and “Bad photo” as the positive and negative prompts, respectively. Following [36], we consider the last convolutional layer of the ResNet50 backbone. Fig. 3(a) shows the result for the positive prompt. We observe that, compared to CLIP-IQA, our model leads to a better alignment with high-quality areas of the image, such as the head of the horse. Similarly, Fig. 3(b) illustrates that QualiCLIP focuses on the most degraded parts of the images when considering the negative prompt, in contrast with CLIP-IQA. The improved alignment between the antonym prompts and the corresponding regions of the images makes QualiCLIP more easily explainable than CLIP-IQA.
S2.5 t-SNE Visualization
@@
S2.6 Supervised QualiCLIP
Although our approach is designed to remove the requirement for human annotations, it can easily be extended to leverage ground-truth labels. Similar to CLIP-IQA+ [47], we exploit the MOS to fine-tune the antonym prompts with an MSE loss via standard backpropagation, while keeping the network weights fixed. We refer to this supervised opinion-aware variant of our approach as QualiCLIP+. We train QualiCLIP+ on the training split of each of the testing datasets and report the performance in Tab. S5. For a fair comparison, we only consider supervised baselines that do not fine-tune the network weights but instead train a smaller set of parameters, such as a linear regressor or the antonym prompts. We observe that QualiCLIP+ achieves competitive performance also in this evaluation setting, outperforming the baselines on most metrics. This outcome shows that the proposed method can also be applied in scenarios where human annotations are available.
Authentic Datasets Image Restoration Datasets AIGC Datasets
KonIQ-10k CLIVE FLIVE SPAQ CVIU PIPAL AGIQA-1K AGIQA-3K
Method OU SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
CONTRIQUE [26] 0.874 0.882 0.806 0.838 0.596 0.648 0.910 0.912 0.912 0.916 0.581 0.611 0.799 0.855 0.817 0.879
Re-IQA [35] 0.883 0.887 0.783 0.797 0.623 0.685 0.909 0.912 0.887 0.895 0.568 0.587 0.783 0.840 0.811 0.874
ARNIQA [1] 0.869 0.883 0.797 0.823 0.595 0.670 0.904 0.909 0.919 0.926 0.634 0.666 0.768 0.849 0.803 0.881
CLIP-IQA+ [47] 0.873 0.890 0.815 0.849 0.602 0.662 0.901 0.904 0.925 0.918 0.552 0.558 0.817 0.854 0.844 0.894
GRepQ [41] 0.882 0.883 0.793 0.813 0.576 0.611 0.902 0.903 0.882 0.857 0.554 0.568 0.740 0.797 0.807 0.858
QualiCLIP+ 0.889 0.898 0.821 0.867 0.618 0.665 0.911 0.914 0.936 0.936 0.586 0.590 0.826 0.871 0.860 0.903
Table S5: Comparison between QualiCLIP+ and supervised approaches. OU stands for Opinion-Unaware. Best and second-best scores methods are highlighted in bold and underlined, respectively.
S2.7 Inference Time
@@
To validate this, we compare the average inference time of our model with that of the supervised baselines on the KonIQ-10k dataset [11], which comprises 10K $1024\times 768$px images. We conduct the experiments on an NVIDIA RTX 2080Ti GPU. We report the results in Tab. S6. As expected, QualiCLIP has a similar inference time to models based solely on ResNet50, such as CLIP-IQA+, LIQE, CONTRIQUE, and ARNIQA. Note that CONTRIQUE and ARNIQA employ the full- and half-scale versions of the input image to compute the final quality score, and thus are more computationally expensive. Moreover, our model is faster than methods based on two encoders, namely GRepQ and Re-IQA. This outcome demonstrates the efficiency and applicability of QualiCLIP in real-world scenarios.
@@
Table S6: Comparison of the average inference time between QualiCLIP and supervised methods on the KonIQ-10k [11] dataset.
@@
Training details We rely on a ResNet50 [10] as the backbone for CLIP’s image encoder. Similar to [47], we remove the positional embedding from the encoder to make our model capable of taking images of any resolution as input. The dimension $d$ of CLIP’s embedding space is 1024. Differently from [41], we do not train a projector head on top of CLIP’s image encoder. We keep CLIP’s text encoder frozen. We train our model for 10 epochs. We employ an AdamW [20] optimizer with a weight decay and a learning rate of $21e\!-\!$ 2 and $91e\!-\!$ 9, respectively. During training, we use a patch size of 224 and a batch size of 16. We set the margins $m_{cons}$ in Eq. 1 and $m_{rank}$ in Eq. 2 and 3 to $32.5e\!-\!$ 3 and $26.75e\!-\!$ 2, respectively. The loss weights $\lambda_{cons}$, $\lambda_{pos}$ and $\lambda_{neg}$ in Eq. 4 are all equal to 1. We set the temperature hyperparameter $\tau$ in Eq. 5 to 2. At inference time, our model takes the whole image as input.
Prompts Following [48, 49], we employ multiple pairs of antonym prompts during training and inference. In particular, we use: 1) “Good/Bad photo”; 2) “Good/Bad picture”; 3) “High-resolution/Low-resolution image”; 4) “High-quality/Low-quality image”; 5) “Sharp/Blurry image”; 6) “Sharp/Blurry edges”; 7) “Noise-free/Noisy image”. We average the similarities between the images and the pairs of prompts.
Synthetic distortions As detailed in Sec. 3.2, during training we synthetically degrade pristine images with increasing intensity levels. Specifically, similar to [1], we consider 24 distinct distortion types divided into the 7 degradation groups defined by the KADID-10k [18] dataset. Each degradation has 5 degrees of progressively higher intensity. We report an example for all the intensity levels of each type of degradation in Figs. S5, S6, S7, S8, S9, S10 and S11. Each distortion is described as follows: